Speed Racer accelerates from rest at a constant rate of 2 m/s2. How fast will the racer be going at the end of 6 s?


1
Expert's answer
2021-09-08T16:14:55-0400
"v=v_0+at,""v=0+2\ \dfrac{m}{s^2}\cdot6\ s=12\ \dfrac{m}{s}."